### v1

# Exploring the Effects of User-Agent and User-Designer Similarity in Virtual Human Design to Promote Mental Health Intentions for College Students

Abstract.
Virtual humans (i.e., embodied conversational agents) have the potential to support college students’ mental health, particularly in Science, Technology, Engineering, and Mathematics (STEM) fields where students are at a heightened risk of mental disorders such as anxiety and depression. A comprehensive understanding of students, considering their cultural characteristics, experiences, and expectations, is crucial for creating timely and effective virtual human interventions. To this end, we conducted a user study with 481 computer science students from a major university in North America, exploring how they co-designed virtual humans to support mental health conversations for students similar to them. Our findings suggest that computer science students who engage in co-design processes of virtual humans tend to create agents that closely resemble them demographically–agent-designer demographic similarity. Key factors influencing virtual human design included age, gender, ethnicity, and the matching between appearance and voice. We also observed that the demographic characteristics of virtual human designers, especially ethnicity and gender, tend to be associated with those of the virtual humans they designed. Finally, we provide insights concerning the impact of user-designer demographic similarity in virtual humans’ effectiveness in promoting mental health conversations when designers’ characteristics are shared explicitly or implicitly. Understanding how virtual humans’ characteristics serve users’ experiences in mental wellness conversations and the similarity-attraction effects between agents, users, and designers may help tailor virtual humans’ design to enhance their acceptance and increase their counseling effectiveness.
embodied conversational agents, virtual humans, mental health, similarity-attraction, participatory design
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TAPjournalvolume: Xjournalnumber: Xarticle: Xccs: Computing methodologies Intelligent agents
1. Introduction
This paper discusses the role of the similarity-attraction effect (Bernier and Scassellati, 2010) when users act as co-designers of virtual humans (i.e., embodied conversational agents), particularly its effects on their perceptions and intentions in sensitive areas like mental health counseling. We also delve into how demographic similarities between users, virtual humans, and designers influence computer science students’ willingness to engage in mental health conversations. Virtual humans are computer-generated characters that resemble humans in appearance, behavior, and ability to communicate using natural language (Guadagno et al., 2007). They have been successfully employed in various scenarios, such as training medical students (Zalake et al., 2021b; Yao et al., 2020) as well as providing mental health support (Parmar et al., 2018; Yao et al., 2020; Ali et al., 2020). In scenarios of mental health support, the lack of rapport between counselors and patients can result in premature patient withdrawal (Hatcher et al., 2005; Zlotnick et al., 1998). Designing virtual humans for mental health support can be challenging, as these scenarios also require the development of rapport between users and virtual humans (Blow et al., 2008; Wintersteen et al., 2005). Therefore, understanding the users’ identities, needs, and expectations when designing virtual humans to assist in mental health contexts is critical (Feijóo-García et al., 2021, 2022, 2023) and can help in increasing rapport among users and agents (Lucas et al., 2014).
Undergraduate college students have been found to struggle with mental health problems such as anxiety (Soares Passos et al., 2020; Gorenstein et al., 1995), depression (Gorenstein et al., 1995), as well as the impostor phenomenon (Rosenstein et al., 2020): i.e., highly-accomplished individuals perceiving themselves as frauds (Clance and Imes, 1978). Similarly, research has reported on graduate students experiencing up to six times higher rates of mental health problems compared to individuals in their same age range (Evans et al., 2018). Within college populations, students who major in Science, Technology, Engineering, and Mathematics (STEM) are at higher risk for mental health problems (Soares Passos et al., 2020). For instance, undergraduate engineering students experience higher rates of anxiety and depression than general undergraduate students (Danowitz and Beddoes, 2018), and among all STEM majors, computer science students have been found to experience the highest rates of anxiety and depression (Danowitz and Beddoes, 2018), with female computer science students being particularly more at risk (Rosenstein et al., 2020; Soares Passos et al., 2020).
Prior research in social psychology has examined the similarity-attraction effect (Bernier and Scassellati, 2010) in counselor-patient matches in mental health contexts, focusing on personality (Byrne et al., 1967; Singh, 1974; Chen and Kenrick, 2002), attitude (Clore and Baldridge, 1968; Condon and Crano, 1988), gender (Bhati, 2014; Zane and Ku, 2014), race (Cabral and Smith, 2011), and ethnicity (Cabral and Smith, 2011; Zane and Ku, 2014). Studies have also explored this effect in virtual human design, considering aspects like gender (Bhati, 2014; Zane and Ku, 2014; Feijóo-García et al., 2023), ethnicity (Baylor and Kim, 2004; Baylor et al., 2003; Feijóo-García et al., 2023), and language (Feijóo-García et al., 2021, 2022, 2023). For example, accent similarity influences computer science students’ perceptions of virtual humans’ expertise (Feijóo-García et al., 2021) and conversational abilities (Feijóo-García et al., 2022). However, research on virtual human design related to agent-user similarity is limited. For instance, as of 2018, only four papers from the Association of Computing Machinery (ACM) International Conference on Intelligent Virtual Agents (IVA) addressed agents’ identities concerning cultural aspects (Norouzi et al., 2018).
Literature on the designer as a product cue has reported its advantages in various markets and products: Customers’ perceptions of products can be shaped by the information they have about their creative actors and the design processes behind them (Da Silva et al., 2015; Fuchs et al., 2015). As a result, having the designer as a product cue could improve customers’ perceptions of a product’s overall quality (Idemen et al., 2021). Nonetheless, research on virtual human design has not yet explored the designer’s role as a product cue and the extent to which virtual humans’ characteristics reflect the creative actors behind their designs–i.e., the designer’s legacy. Also, although the designer as a product cue has been used in recent years as an extrinsic cue to promote products in various markets (e.g., ”designed by”) (Idemen et al., 2021), it remains unclear how extrinsic this cue may be in the context of virtual human design and to what degree end users expect virtual humans to resemble their designers. Therefore, this paper reports on the influence of the designer’s legacy (i.e., designer-agent demographic similarity) on virtual humans created to encourage computer science students’ intentions toward mental health conversations. Our work explored how computer science students perceived designers’ legacies (i.e., designer-agent demographic similarity) on virtual humans, when asked to describe the designers behind the virtual humans they interact with. We aimed to determine how designers’ characteristics were perceived and manifested through the virtual human as an interactive interface, and to understand the similarity-attraction effect (Bernier and Scassellati, 2010) by examining whether students’ motivation and interest in a virtual human, when students acted as co-designers, increased when the agent’s designer was demographically more similar to the end user student.
Our research builds upon the research conducted by Feijóo-García et al. (Feijóo-García et al., 2023), who explored how computer science students co-designed virtual humans to support mental wellness conversations for target users similar to them. Co-design has proven effective in enhancing user adoption (Frauenberger et al., 2011; Muller and Kuhn, 1993; Van Mechelen et al., 2021), improving user input (Wobbrock et al., 2009; Van Mechelen et al., 2021), and empowering user communities by giving them more control over solutions that impact their lives (Frauenberger et al., 2011; Vasalou et al., 2021; Simonsen and Robertson, 2013). Feijóo-García et al. (Feijóo-García et al., 2023) report on students leaning toward user-agent (or agent-designer in this case) demographic similarity in regard to gender, age, ethnicity, and their places of origin when they participate in virtual human co-design. Moreover, the authors provided insights on how the interplay between the virtual human’s appearance and voice impacted computer science students’ intentions toward gratitude journaling (Feijóo-García et al., 2023).
The findings by Feijóo-García et al. (Feijóo-García et al., 2023) contributed to the understanding of the similarity-attraction effect in co-design scenarios of virtual humans, especially in regard to user-agent demographic similarity. However, their work had limitations concerning the size of their sample population (their first user study had a total of 73 computer science students acting as co-designers of virtual agents) and reported missing opportunities concerning the exploration of user rationale in the selection of virtual humans’ appearances and voices, as well as how both characteristics interplay concerned designers’ demographic similarity expectations (Feijóo-García et al., 2023). Therefore, continuing their efforts, our user study went beyond and far extended their research with the aim of better understanding how computer science students co-designed virtual humans to support. Our research had a larger and more representative sample population (481 computer science students) and extended the protocol used by Feijóo-García et al. (Feijóo-García et al., 2023) to include participants’ rationale behind the selection of the agents’ demographic characteristics, as well as their appearances and voices, to analyze the interplay between appearance, voice, and agents’ desired and perceived demographics.
The research we report in this paper involved a two-part user study that extended the work by Feijóo-García et al. (Feijóo-García et al., 2023). The first part of this user study investigated how virtual humans designed by computer science students were influenced by demographic similarity expectations–students willing their agents to be like them, reporting insights on the interplay among virtual humans’ characteristics, appearances, and voices, when co-designed by computer science students to support someone similar to them in their mental health. The second part of this user study explored whether end users’ attraction to virtual humans increased with a designer more demographically similar to them when designers’ characteristics were implicitly or explicitly used as product cues. Additionally, the second part of our user study explored how end users perceived designers’ demographic characteristics through their agents, reporting insights on end users’ expectations of agent-designer demographic similarity.
Our research aimed to respond to the following research questions:
  • RQ-“Similarity Elucidation”: How do similarity expectations influence the design decisions made by computer science students when co-designing virtual humans for mental health support?
  • RQ-“Designer Cues”: Which characteristics of a virtual human’s designer can be identified by computer science students during their interactions with the agent?
  • RQ-“Designer’s Legacy Effectiveness”: How effective are virtual humans in promoting computer science students’ intentions toward mental wellness conversations when co-designed by computer science students with similar demographic characteristics to their end users?
2. Related Literature
2.1. The Designer as a Product Cue
Prior research in marketing has reported the impacts of having the designer as a product cue. That is, to allow end consumers to know who is behind the products they use (Idemen et al., 2021). Studies have reported on how customers’ perceptions of products are shaped by the information they have about their creative actors and the design processes behind them (Da Silva et al., 2015; Fuchs et al., 2015). Fuchs et al. (Fuchs et al., 2015) conducted a set of studies that compared handmade and machine-made products, evaluating how the stated mode of production impacted products’ attractiveness. The authors found that consumers indicated stronger purchase intentions and paid more for handmade products when buying gifts for their loved ones, suggesting that customers had a special appreciation for the human factor in production. Similarly, Da Silva et al. (Da Silva et al., 2015) conducted a set of studies to evaluate how the knowledge of designers’ intentions (or rationale) impacted users’ appreciation of products. The authors argued that a product could present designers’ intentions through the product itself or via marketing and advertisement. Their findings suggest customers’ appreciation of a product can increase when they know the designers’ intentions. However, customers’ appreciation may decrease when the revealed intentions mismatch customers’ initial perceptions of the product. Designers must identify and decide what intentions should be shared and how to the end users.
Having the designer as a product cue can increase how customers perceive the overall quality of a product (Idemen et al., 2021). Idemen et al. (Idemen et al., 2021) conducted a study to identify customers’ insights concerning the use of the designer as a product cue. Their work was interview-based and used a discovery-oriented grounded theory approach. The authors’ found a designer cue can affect how customers perceive a product’s quality, as information about the designer and the design process can reveal the effort and intellectual background behind a product. The authors present that designer cues can also signal product quality and support customers’ reflective level of the experiences they expect with the products. Likewise, their findings suggest designer cues can signal a product’s social value and status and expose the design process and involvement behind a product.
Research has found customers care more about the creative actors behind products in fashion (Joy et al., 2014), music (Smith et al., 2016), cinema (Hennig-Thurau et al., 2007), art (Chen, 2009; Moulard et al., 2014), and wine-making industries (Beverland, 2005). Wu et al. (Wu et al., 2017) found customers evaluate products more positively when they know about the process and effort invested in their designs (Wu et al., 2017). The authors also found that customers appreciated highly aesthetic products more, as these products indicated more effort in their design–these findings are similar to those by Idemen et al. (Idemen et al., 2021). Nevertheless, research on virtual human design has yet to explore the role of the designer as a product cue, as literature about this topic is still missing. Also, research is needed to understand how virtual humans’ characteristics reflect the creative actors behind their designs–i.e., the designer’s legacy, and the impact of the designer’s legacy on users who interact with virtual humans in sensitive contexts such as mental health.
2.2. User Participation in Virtual Human Design
Several studies on virtual human design have adopted diverse methods in shaping the conversational design of agents. In the context of virtual human patients, approaches have gone from using persona-based design (Whittaker et al., 2021) to characterizing agents’ dialogues based on their narrative, the problems they aim to address (Bearman et al., 2001), as well as promoting agents’ credibility by using corpus-based design for question-answering scenarios (Halan et al., 2014).
Previous research has also actively involved users in the design of virtual humans. Halan et al. (Halan et al., 2015) investigated how racial similarity impacted the design of virtual human patients in the scenario of promoting empathetic abilities among medical trainees. They had participants design and interview the virtual human patients. Remarkably, they found a notable difference in empathy ratings between groups with user-agent racial concordance and discordance when the virtual patient was Afro/Black American and the user Caucacian, suggesting that the effect of creating a virtual human patient could be more pronounced when there is user-agent racial discordance. Similarly, Ghosh et al. (Ghosh et al., 2023) explored how gender is modeled and designed in virtual humans with computer science students engaging in their co-design process. Their findings suggest that gender cues in virtual human faces, such as age, hair, makeup, and facial features, are often stereotypically applied for binary genders (male and female) and inconsistently used for non-binary virtual humans. This inconsistency suggests a lack of clear guidelines or norms for representing non-binary individuals in virtual human design, highlighting the need for more inclusive and varied representations of gender in virtual environments (Ghosh et al., 2023).
A recent study by Feijóo-García et al. (Feijóo-García et al., 2023) had computer science students co-design virtual humans to assist students similar to them in their mental well-being. Their research found that computer science students leaned toward creating agents closely resembling their own demographic characteristics, with Black, Indigenous, and People of Color (BIPOC) participants preferring user-agent ethnic similarity and female participants preferring user-agent gender similarity. Age was also found critical, as it was found that participants negatively perceived a virtual human whose age exceeded their expected maximum age. Additionally, Feijóo-García et al. reported on how the interplay between a virtual agent’s appearance and voice could impact the agent’s effectiveness in promoting computer science students’ intentions toward gratitude journaling (Feijóo-García et al., 2023). These findings relate to those by Yin et al. (Yin et al., 2010), who assessed the interplay between virtual human appearance (”Latin American”–dark-skinned or ”Anglo-American”–light-skinned) and language (English or Spanish) and its effectiveness on user persuasion. As they report, ”Many subjects mentioned that the agent’s language strongly influenced their perception of the agent’s cultural background” (Yin et al., 2010).
Our user study builds upon the foundational work of Feijóo-García et al. (Feijóo-García et al., 2023), delving into the dynamics of user-agent similarity in the design of virtual humans, particularly in scenarios where users partake in the co-creation process–agent-designer demographic similarity in this case. By investigating how computer science students co-design virtual agents that reflect aspects of their own demographics - such as gender, ethnicity, and age - our research contributes to the existing body of knowledge on the application of virtual humans in mental health support scenarios, particularly in regard college students needs. Our research not only echoes the findings of Feijóo-García et al. (Feijóo-García et al., 2023) about the tendency of designers to create agents resembling themselves but also amplifies the understanding of how these virtual entities can be optimally designed to foster mental wellness conversations among college student populations.
2.3. The Similarity-Attraction Effect on Virtual Human Design
Previous research on virtual human design has used different angles to understand the impact characteristics such as gender, culture, and language have on users’ interactions with computer systems such as virtual humans.
2.3.1. Gender similarity
Research in human-computer interaction reveals that gender significantly shapes user preferences towards interactive systems like robots and virtual humans. Studies, such as those by Kulms et al. (Kulms et al., 2011), and Stal et al. (ter Stal et al., 2020), often show that female virtual agents are viewed more positively than male agents. However, the picture is more complex, as research by Guadagno et al. (Guadagno et al., 2011), and Kim and Baylor (Kim et al., 2007) indicates a general preference for virtual humans of the same gender as the user, with perceptions varying based on both the user’s and the agent’s gender. Despite some positive views of female agents, Baylor et al. (Baylor and Kim, 2004) observed that they are often perceived as less knowledgeable than male agents. Regarding gender recognition, Nag and Yalçın (Nag and Yalçın, 2020) found a tendency to categorize androgynous virtual agents as male.
Expanding on these findings, the idea of user-agent similarity-attraction based on gender is clearly visible in this field. Lee et al. (Lee et al., 2000) reported a preference among users for systems that match their own gender, though male voices were generally favored, reflecting societal stereotypes. This intersection of culture and gender is further highlighted in Koda and Takeda’s (Koda and Takeda, 2018) study, where both gender and cultural gaze behaviors were found to influence Japanese participants’ views on virtual agents. Finally, Baylor et al. (Baylor et al., 2003) noted in an educational context that male students preferred male pedagogical agents, emphasizing gender’s significant influence in virtual learning environments.
2.3.2. Race and ethnicity similarity
Research has extensively explored the roles of race, culture, and ethnicity in human-computer interaction, particularly in the design of virtual agents and interfaces (Rossen et al., 2008; Pratt et al., 2007; Baylor and Kim, 2004; Iacobelli and Cassell, 2007; Baylor et al., 2003; Wang et al., 2020; Jan et al., 2007; Lucas et al., 2018; Koda and Takeda, 2018). Koda and Takeda (Koda and Takeda, 2018) found that Japanese participants’ impressions of virtual agents varied based on cultural familiarity with the agents’ gaze behaviors. Lucas et al. (Lucas et al., 2018) reported that cultural background (Japanese or American) influenced rapport-building with agents, particularly in responses to ice-breakers and dialogue errors. Wang et al. (Wang et al., 2020), applying Hofstede’s six-dimension model (Hofstede, 2010), observed cultural differences in reactions to agents’ feedback, with American users showing more resistance to negative feedback.
User studies on race and ethnicity in educational contexts (Baylor and Kim, 2004; Baylor et al., 2003) revealed that the agents’ ethnicity influenced participants’ learning outcomes. For instance, Baylor and Yanghee (Baylor and Kim, 2004) found better learning transfer with Black (i.e., dark-skinned) virtual humans compared to White (i.e., light-skinned or Caucasian) ones. Iacobelli and Cassell (Iacobelli and Cassell, 2007) observed that ethnicity recognition in virtual agents could occur through language use, as African American children identified and preferred agents speaking African American Vernacular English. Pratt et al. (Pratt et al., 2007) and Baylor et al. (Baylor et al., 2003) also found that students preferred agents of their own ethnicity, indicating a similarity-attraction effect.
However, despite these insights, only a few studies have directly addressed the similarity-attraction effect and its impact on user interaction with culturally relevant virtual humans (Iacobelli and Cassell, 2007; Pratt et al., 2007; Baylor et al., 2003). Further research is necessary to fully comprehend this aspect of human-computer interaction, particularly for virtual human design.
2.3.3. Language similarity
Language is critical in human-computer interaction (HCI), influencing user preferences and perceptions in various domains. In social robotics, language plays a crucial role in creating culturally relevant systems (Lugrin et al., 2018; Oliveira et al., 2020; Joosse et al., 2013), with humor modeling (Oliveira et al., 2020) enhancing users’ positive perceptions. Cultural adaptation, based on models like Hofstede’s (Hofstede, 2010), is also significant in human-robot interaction (HRI) (Lugrin et al., 2018). Additionally, studies show users prefer robots with similar personalities to their own (Joosse et al., 2013).
For speech assistants, research indicates a preference for human voices over synthetic ones (Cambre et al., 2020), with multiple factors like gender, age, accent, and vocal characteristics influencing user interaction (Seaborn et al., 2021). Technical limitations still exist in adequately addressing varied speech patterns (Sutton et al., 2019). Parviainen and Søndergaard (Parviainen and Søndergaard, 2020) examined how whispering affects voice assistant experiences, highlighting the importance of voice tone and proxemics. Braun et al. (Braun et al., 2019) found personalized voice assistants enhance trust, particularly when matching the user’s personality.
In virtual human research, Iacobelli and Cassell (Iacobelli and Cassell, 2007) used language to model non-visual ethnicity, demonstrating that African American participants could identify agent ethnicity based on language use, with a preference for agents that matched their own ethnic background. Studies on spoken interfaces like Dahlbäck et al. (Dahlbäck et al., 2007) reveal a similarity-attraction effect, with users preferring accents similar to their own.
Despite these findings, there’s limited research on the impact of language and its properties, such as accent, on virtual human design, particularly concerning the similarity-attraction effect. Our work extends the work by Feijóo-García et al. (Feijóo-García et al., 2023) and aims to explore language’s role in user engagement with virtual humans, particularly in sensitive contexts like mental health. This paper reports on how users as co-designers of virtual humans influence perceptions and intentions toward coping techniques such as gratitude journaling (Flinchbaugh et al., 2012).
2.4. Counselor-Patient Matching in Mental Health Scenarios
The principle of similarity-attraction states that people prefer to interact with others they perceive as similar to themselves (Bernier and Scassellati, 2010). Previous research in social psychology has reported on the effects of similarity-attraction concerning mental health setups among patients and therapists, considering multiple factors such as 1) personality (Byrne et al., 1967; Singh, 1974; Chen and Kenrick, 2002), 2) attitude (Clore and Baldridge, 1968; Condon and Crano, 1988), 3) gender (Bhati, 2014; Zane and Ku, 2014), 3) race (Cabral and Smith, 2011), and 4) ethnicity (Cabral and Smith, 2011; Zane and Ku, 2014).
2.4.1. Counselor-patient ethnic matching
Cabral and Smith (Cabral and Smith, 2011) conducted a meta-analysis of over 80 references about patients’ preferences towards counselors regarding their gender and ethnicity. They reported that the enhancement (i.e., increasing the number of common characteristics) of counselor-patient similarity had no impact on mental health treatment. Treatment outcomes were not found to be different when patients experienced a process with a therapist who shared their same race or ethnicity. However, the authors argued that it could be due to biases according to interpersonal-similarity beliefs: an assumption of more remarkable similarity than the actual one. Their analysis suggests that patients preferred having a therapist with the same race and ethnicity: racial and ethnic matching could initiate counseling. The authors found that once patients had initiated the therapeutic treatment, race, and ethnicity became secondary and did not generally impact how much benefit they received. However, the authors also found interesting insights concerning African Americans. Different from other ethnic and racial groups, African American patients strongly preferred African American counselors.
2.4.2. Counselor-patient gender matching
Concerning gender and ethnicity with Asian participants, Zane and Ku (Zane and Ku, 2014) studied whether ethnic and gender similarity impacted how patients self-disclosed (about feelings, habits, and sexual issues) with counselors. Their population consisted of 110 Asian American students from a North-American college. The authors did subgroup analysis on different Asian ethnic subgroups: characterized as Chinese Americans (45%), Japanese Americans (18%), Korean Americans (26%), and Vietnamese Americans (11%). The authors’ methodology asked participants to listen to an audio recording of a counselor, which could display the characteristics of an Asian male or female or a White (i.e., Caucasian or of European origin) male or female. The study reports a similar position to Cabral and Smith (Cabral and Smith, 2011): Ethnic matching did not affect patients’ self-disclosure. However, gender matching led patients to self-disclose more.
Gender was also explored by Bathi (Bhati, 2014), who followed a naturalistic (or in-situ) methodology and hypothesized gender to be meaningless towards therapeutic alliance once treatment was initiated. Their study assessed 92 adults diagnosed with a mental disorder. Participants were primarily White Americans (93.5%) and females (76.1%). Participants were assigned to counselors that matched or did not their gender: 49 pairs matched, and 43 did not. The study (Bhati, 2014) was conducted for 15 weeks, and data was collected every three to five sessions. Gender similarity was found to impact therapeutic alliance between patients and counselors positively. Moreover, the authors found that female patients matched with female counselors reported higher therapeutic alliance ratings than male counselor-patient pairs.
2.4.3. Counselor-patient attitude matching
Social psychologists have also explored the similarity-attraction effect concerning attitude and interpersonal agreement (Clore and Baldridge, 1968; Condon and Crano, 1988; Chen and Kenrick, 2002). Clore and Baldridge (Clore and Baldridge, 1968) reported results from one study designed to evaluate interpersonal attraction. Their study assessed 84 participants from a North-American college who were asked to respond to a 56-item attitude survey that included 4-point scales that measured participants’ interest per item. Then, the researchers randomly distributed participants’ anonymously answered surveys, asking them to judge their peers’ responses based on criteria such as intelligence, knowledge, morality, and how likely they would like to pair with their matching peers. The authors found that participants could be attracted to a stranger when they happened to agree more (Clore and Baldridge, 1968).
Similarly, Condon and Crano (Condon and Crano, 1988) evaluated how attitude-attraction between individuals varied on how they assessed others. They recruited 260 first-year undergraduate students from a North American college. They asked them to answer a questionnaire that collected their ratings on religion, political party, premarital sex, etc. Then, the authors randomly assigned participants’ anonymously responded surveys among them, asking participants to judge their peers’ interests. As Clore and Baldrige (Clore and Baldridge, 1968) report, Condon and Crano (Condon and Crano, 1988) found Agreement to affect how individuals felt attracted to each other —these findings can contribute to the design of virtual humans that provide feedback to users (De Visser et al., 2016; Jin and Youn, 2021). Fang and Kenrick (Chen and Kenrick, 2002) evaluated similarity-attraction and dissimilarity-repulsion concerning group membership—e.g., belonging to the same political party. Two of their studies considered attraction and repulsion on 1) political affiliation (Democrat or Republican) and 2) sexual orientation (Heterosexual or Homosexual). The authors found that participants experienced a higher attraction towards groups they shared similarities with. Additionally, they found that groups that participants knew were different fostered higher participants’ repulsion.
2.4.4. Counselor-patient personality matching
Personality has also been considered when evaluating the similarity-attraction effect in counselor-patient setups (Byrne et al., 1967; Singh, 1974). Byrne et al. (Byrne et al., 1967) assessed 450 participants—male and female undergraduates from a North American college, asking them to respond to the Repression-Sensitization (R-S) Scale (Byrne et al., 1963) to measure their personality. As done in studies for attitude-similarity (Clore and Baldridge, 1968; Condon and Crano, 1988), Byrne et al. (Byrne et al., 1967) asked participants to respond to a survey that collected their interests on specific topics. Participants then had to judge their peers’ responses, agreeing or disagreeing with them. Their results suggested that personality correlated with attitude, as participants were more attracted to individuals with similar personality attitudes. Similarly, Sing (Singh, 1974) evaluated the role of affective states concerning similarity-attraction, assessing 143 male and 127 female undergraduate first-year students from a North American college. Their personality was measured with the scale proposed by Byrne (Byrne et al., 1963), and participants also had to judge their peers based on the limited information provided by an already filled survey (Singh, 1974). The authors’ findings also suggest that personality-similarity increases the attraction between individuals, as also attitude-similarity does (Singh, 1974).
3. Study Design, Part 1: Co-Design of Virtual Humans for Mental Health Support
The user study took place in the Spring of 2023. It was conducted online and asynchronously, and consisted of two distinct parts. Both parts were separated by a period of over four days, allowing sufficient time to create tailored virtual humans for the second part of the study. The user study was approved by the Institutional Review Board (IRB) prior to its execution (Protocol #17817).
The first part of this user study helped explore designer-agent demographic similarity when computer science students engaged in virtual human co-design intending to design virtual humans to assist toward mental health scenarios (Figure 1). This first part addressed the following research question:
RQ-“Similarity Elucidation”: How do similarity expectations influence the design decisions made by computer science students when co-designing virtual humans for mental health support?
Figure 1. Two-Part User Study Overview: User-Designer-Agent Demographic Similarity in Mental Health
3.1. Participants
The first part of this user study had a sample population that consisted of adult graduate and undergraduate computer science students from the University of Florida. A total of 481 adult (18 years of age or older) participants were recruited regardless of gender, age, language, ethnicity, or place of origin. Participants were enrolled in the Spring of 2023 and came from a diverse range of courses: two graduate-level Human-Centered Computing courses, as well as six undergraduate computer science courses, including two freshman, one sophomore, and three senior computer science courses.
Table 1. Ethnic Groups per Gender
Ethnic Groups per Gender
Ethnic Group Total Gender
Female Male Other
Asian n=217 n=94 n=118 n=5
Afro/Black American n=28 n=11 n=15 n=2
Latin American n=100 n=38 n=57 n=5
Middle-Eastern/North-African n=14 n=4 n=9 n=1
Native American/Alaska Native n=2 n=1 n=1 n=0
Pacific Islander n=8 n=3 n=4 n=1
White (non-Latin American) n=202 n=63 n=126 n=13
Participants’ reported ages were: 18-20 years of age (n=339), 21-25 years of age (n=119), 26-30 years of age (n=8), and over 30 years of age (n=13). Two participants (n=2) did not disclose their age. Participants also reported the number of languages they were familiar with–monolingual (n=197) or multilingual (n=284), and their place of origin–country and city. Participants reported being from 34 different countries: from the United States (n=355) or other places (n=126)–see Appendix A, Table 9. Participants also reported being domestic (n=402) or international (n=79) students.
Participants were also asked to report their gender and ethnicity. The gender distribution among participants included Female (n=182), Male (n=281), and Other (n=18)—e.g., gender-queer, non-binary, transgender. Table 1 presents the participants’ ethnic groups categorized by their gender. Each row should be considered independently, as there was a group of multi-ethnic participants (n=76) who selected more than one ethnic group.
Figure 2. Part 1: Procedure
3.2. Procedure
The first part of the user study was conducted online and asynchronously, and took up to 30 minutes for participants to complete it. Participants were asked to respond to an online questionnaire, in which they were tasked with designing a mental health virtual human to assist someone ”similar to them”. Figure 2 illustrates the procedure of the first part of this user study. Questions were organized into seven blocks:
Block 1-Informed Consent provided an informed consent document with details of the study and two multiple-choice questions. Participants were asked to read it and respond 1) if they agreed to participate and 2) if they were 18 years of age or older. Participants were informed and agreed to share institutional emails, instructors’ names, and courses’ names.
Block 2-Participant’s Information included three questions that asked participants to input their institutional email, institutional emails, instructors’ names, and courses’ names.
Block 3-Virtual Human’s Demographics asked participants to describe the virtual human they wanted to design. The block included four questions concerning the virtual human’s demographic characteristics: age (open-ended numeric input), ethnicity (multiple-choice close-ended question based on (Feijóo-García et al., 2021)), place of origin (open-ended input as country-city), and gender (close-ended multiple-choice question based on (CSUSM, 2022)).
Block 4-Virtual Human’s Appearance and Voice focused on the appearance of the virtual human, asking participants to respond to six questions to select a male and a female appearance. The catalog of virtual humans featured 12 male and 15 female appearances from agents available in Synthesia (Synthesia, 2022) that could be used with the Synthesia API (syn, 2023). The first three questions focused on selecting a male appearance. The first question asked participants to select the top three male appearances they considered would encourage students similar to them to engage in mental wellness conversations. The second question asked participants to select the voice considered most suited from the previous three selected voices for the top male they decided before. Finally, the last question asked them to explain their choice. The last three questions of this block followed the same format to select a female appearance.
Block 5-Participant’s Perceptions of the Virtual Humans’ Demographics presented four close-ended questions. Participants were first asked to indicate the perceived age (open-ended numeric input) and ethnicity (multiple-choice close-ended question based on (Feijóo-García et al., 2021)) for the top male appearance they selected. Next, they did so for the top female appearance they chose.
Block 6-Virtual Human’s Voice centered on the virtual human’s voice and presented six questions for participants to select a male and a female voice. The catalog of voices featured 12 male and 12 female voices available in Synthesia (Synthesia, 2022) that could be used with the Synthesia API (syn, 2023)–see Table 15 for male voices, and Table 16- for female voices. All voices featured a five-second utterance (to limit mental load) and were labeled with their nationality (e.g., Canada). The first three questions focused on selecting a female voice. The first question asked participants to select the top four female voices they considered would encourage students similar to them to engage in mental wellness conversations. The second question asked them to select the top choice from the previous four selected options. Finally, the last question asked them to explain their top choice. The last three questions of this block followed the same format to select a male voice.
Block 7-Participants’ Demographics focused on participants’ demographics, presenting six questions. First, it asked about participants’ age (open-ended numeric input), ethnicity (multiple-choice close-ended question based on (Feijóo-García et al., 2021)), and gender (close-ended multiple-choice question based on (CSUSM, 2022)). It followed by asking about participants’ international student status (USA domestic or international) and their place of origin (open-ended input as country-city). Finally, it asked one open-ended question to have participants list all languages used to communicate.
3.3. Data analysis
Data from the first part of this user study was analyzed with a focus on agent-designer demographic similarity, comparing the characteristics of the virtual humans selected by the participants to the participants’ own demographic characteristics reported after participating in the first part of the user study. The analysis considered gender, ethnicity, age (with a similarity window of ±5 years (Feijóo-García et al., 2023)), and place of origin. We used the Chi-Squared Test categorical data to evaluate the statistical association between 1) participants’ self-reported genders and the gender preferences for the virtual agents and 2) participants’ self-reported places of origin and the places of origin preferred for the virtual agents concerning their countries. We used the Chi-Squared Test on these categorical responses as they corresponded to questions in which participants could only select one answer (Franke et al., 2012).
The Pearson correlation coefficient was used to analyze user-agent age similarity, assessing the relationship between participants’ self-reported ages and the preferred ages of their virtual agents. We used this statistical method on these responses as they were numerical and corresponded to questions in which participants could only select one answer (Cohen et al., 2009).
We also analyzed the relationship between participants’ self-reported ethnic groups and those ethnicities selected for their virtual agents by calculating the Jaccard similarity coefficient to assess the similarity among response groups. We decided to use this method due to the multi-label format of participants’ responses, as participants could select more than one ethnic group for themselves, as well as multiple ethnic groups for their virtual agents (Luaces et al., 2012).
Moreover, the study examined how participants selected virtual humans’ appearances and voices when prompted to choose options that would encourage ”someone similar to them” to engage in mental wellness conversations. Top preferences on appearances and voices were analyzed using descriptive statistics, cross-referenced with participants’ demographics. We used the Pearson correlation coefficient to compare participants’ self-reported ages with the perceived ages of the agents’ appearances they selected. Additionally, we calculated the Jaccard similarity coefficient to evaluate the relationship between participants’ self-reported ethnic groups and the ethnic groups they perceived the agents’ appearances they selected represented.
Finally, multi-label thematic analysis was employed to collect and examine participants’ insights from two open-ended questions regarding their design rationale for selecting appearances and voices. Cohen’s Kappa (Gwet, 2008; Viera et al., 2005) and the Jaccard similarity coefficient were used to assess the inter-rater reliability of the themes that came from the qualitative analysis (Luaces et al., 2012).
3.4. Findings: Agent-designer matching of demographic characteristics
Of the four characteristics matched between participants and virtual humans they co-designed, age was the only demographic characteristic for which agent-participant (i.e., agent-designer) matching was not the preferred choice among most participants when designing a virtual human to provide mental health assistance to someone similar to them. Three participants (n=3) did not report a counselor’s age. Among the remaining participants who self-reported their age (n=476), 66.9% of participants preferred an older virtual human (more than five years older), 0.4% preferred a younger virtual human (less than five years younger), and 32.8% designed a virtual human as old as they reported to be (having a window of ±5 years of age for similarity): virtual humans’ average age was of 28.6 years of age, reporting a maximum of 60 years of age. The Pearson correlation coefficient revealed a significant association between participants’ self-reported ages and the virtual agents’ ($r=0.257,n=476,p<0.01$), suggesting a positive relationship between them. This correlation suggests that as the participant’s age increases, the agent’s age also tends to increase, which supports the existence of user-agent age similarity when participants design agents to support individuals similar to them in mental well-being.
For the other three demographic characteristics, most participants (n=481) made design decisions that matched their demographics with those of the agents they designed. Ethnic groups were alike among participants and their virtual agents, with a Jaccard index of 0.62 suggesting that more than 60% of the ethnic categories were shared between them (Luaces et al., 2012). These observations highlight potential cultural influences on participants’ preferences based on expectations of ethnic similarity. Likewise, the Chi-squared test revealed a significant association between participants’ self-reported genders and virtual agents’ genders ($\chi^{2}=293.4,df=4,p<0.01$), as well as participants’ self-reported countries of origin and virtual agents’ countries of origin ($\chi^{2}=5839.7,df=1054,p<0.01$).
From our sample population (n=481), 61.1% of participants decided on ethnic matching and 76.9% on matching their country of origin: 24.1% indicated the same city of origin. The sample population had male (n=281), female (n=182), and participants from other gender groups (n=14)–e.g., non-binary, transgender, etc. From the general population, 53.0% of them preferred a female agent, 36.0% a male one, and 9.8% did so for an agent with another gender identity. See Appendix A: Table 10 shows preferences on gender concerning participants’ ethnic groups.
Looking at gender similarity, 72.8% of the general population decided on gender matching. Female participants (n=182) highly preferred a female agent: 90.1% of the female population. Similarly, 78.6% of participants who identified with other gender identities (n=14–e.g., non-binary, gender-queer, transgender) preferred an agent who gender-matched them. Male participants (n=281) also leaned toward a gender-matching agent, although not as pronounced as participants from other gender groups (61.9% of males). Therefore, gender similarity is critical in the design of virtual humans to support non-Male minorities in mental health.
These findings add to those from Feijóo-García et al., (Feijóo-García et al., 2023), indicating that when co-designing virtual humans to provide mental health assistance to someone similar to themselves, participants generally choose demographic characteristics for the virtual humans that match their own.
3.5. Findings: Participants’ rationale on virtual human appearance selection
Participants (n=481) were asked to provide information about their design decisions concerning virtual human appearance selections. Table 2 displays the ten themes identified through the thematic analysis conducted on participants’ criteria.
Table 2. Themes for Appearance Selection
Themes for Appearance Selection
ID Theme Description
A1 Age The participant considered age as a selection criterion.
A2 Approachability
The participant expressed their choice was due to
the agent’s friendly, welcoming, and/or non-judgmental attitude.
A3 Attire The participant referred to the attire of the agent.
A4 Ethnicity The participant considered ethnicity or race for their selection.
A5 Expertise
The participant referred to the perceived knowledge,
experience, or skill of the agent.
A6 Gender The participant considered gender as a selection criterion.
A7 Non-verbal Language
The participant referred to the facial expression
and/or body language of the agent.
A8 Random
The participant did not elaborate on their
rationale or provided a nonsensical response.
A9 Self-similarity
The participant expressed their choice was
based on finding the agent similar to themselves.
A10 Similarity to Trusted Agent
The participant’s choice was based on a former
relationship and trust with a previous
familiar actor (e.g., a counselor, a family member).
Multiple labels were assigned to participants’ responses based on the themes presented in Table 2. Inter-rater reliability was assessed separately for male and female appearance selections. To achieve this, two independent researchers coded 20% of the responses (Gwet, 2008). Following this, Cohen’s kappa (Gwet, 2008) (for each theme) and Jaccard indexes (Luaces et al., 2012) were calculated for both appearance selections: male and female–see Table 11 in Appendix A for inter-rater reliability scores for male and female appearance selections. Inter-rater reliability was high for each theme, with Cohen’s kappa scores exceeding 0.80 (Viera et al., 2005). Additionally, multi-label inter-rater reliability was also high among raters, with Jaccard indexes surpassing 0.70 in both cases (Luaces et al., 2012).
3.6. Findings: Preferences on encouraging virtual human male appearances
Participants (n=481) were asked to choose up to three male appearances based on their potential to encourage someone to engage in mental wellness conversations. Subsequently, they selected their top male appearance and answered two questions regarding their perceptions of the agent’s age and ethnicity. Figure 3 displays the four male appearances most frequently included in participants’ top three selections, as well as the four least selected male appearances. Figure 4 presents the appearances that participants most often chose as their top choice. The four appearances from Figure 3 were also the most preferred top choices, as shown in Figure 4.
Figure 3. Most and Least Selected Male Appearances
Table 3. Participants’ Perceptions of Ethnicity and Age for Their Top Male Appearances
Perceived Age Perceived Ethnicity
Agent
Average
[years of age]
Standard Deviation
[years of age]
Group
% of responses
out of n
M1 (n=115) 30.3 3.6 Asian 96.5%
M2 (n=70) 38.2 5.2 White (non-Latin American) 94.3%
M3 (n=70) 25.5 3.0 White (non-Latin American) 92.9%
M4 (n=67) 31.7 3.9 Asian 98.5%
3.6.1. Perceptions of age and ethnicity of virtual human male appearances
Using the Pearson correlation coefficient, we found that participants’ self-reported ages (n=479, as two participants did not report their age) significantly correlated with the ages of the male agents they selected as their top-one choice ($r=0.116,n=479,p<0.05$). This positive correlation suggests the presence of user-agent age similarity when participants selected the appearances of their male virtual agents.
Participants’ perceptions of the four top male appearances’ ages and ethnic groups are shown in Table 3. As observed in Table 3, the top agents were between 25 and 40 years of age, with the top two (M1 and M3) close to the general average age of 28.6 years previously reported (Section 3.4). Concerning the ethnicity of the male appearances, two of the top four male appearances (Figure 4) were perceived as White (non-Latin American)–M2 and M3, while the other two (M1 and M4) were considered Asian.
Ethnic groups were moderately alike among participants and their male virtual agents, with a Jaccard index of 0.44 suggesting that more than 40% of the ethnic categories were shared between them (Luaces et al., 2012). Upon examining the ethnic groups more closely, it was observed that participants’ preferences leaned towards ethnic similarity. For example, among the Afro/Black American population (n=28), 75.0% of them chose a top appearance they perceived as non-White, with 67.9% selecting a male appearance perceived as Afro/Black American. Similarly, 53.5% of the White participants (n=202) chose an appearance they perceived as White (non-Latin American), 55.3% of the Asian participants (n=217) favored an Asian-looking male appearance, and 62.5% of the Pacific Islander participants (n=8) selected an ethnically matching appearance. Additionally–although not to be generalized due to the small sample population, both Native American/Alaska Native participants (n=2) chose a male appearance they perceived as matching their ethnicity.
These observations suggest potential cultural influences on participants’ preferences based on expectations of ethnic similarity. However, it’s important to note that the Jaccard index concerning male appearances’ selection is 20% below the one reported in Section 3.4. We consider this may be due to a limited catalog of male appearances. We encourage further research to explore options that provide an open canvas to help elucidate and observe participants’ ethnic expectations when designing virtual humans to support individuals similar to them.
Figure 4. Top Preferences of Male Appearances
Of the seven ethnic groups considered, two of them did not tend toward ethnic similarity concerning male appearances. Among Latin American participants (n=100), only 33.0% of them selected a male appearance they perceived as Latin American. Likewise, only 21.4% of Middle Eastern/North African participants (n=14) leaned toward ethnically similar male appearances. As 93.0% of the Latin American and 92.3% of the Middle Eastern/North African population were domestic, further research is needed to understand if this phenomenon is due to limitations on the catalog of appearances or familiarity aspects regarding the under-representation of these ethnic groups in counseling scenarios in the U.S.A.–counselors in the U.S.A. are predominantly White (non-Latin American) (USA, 2017).
3.6.2. Participants’ rationale on male appearance selection
Participants’ preferences on male appearances were highly led by three of the ten themes used to label data (Section 3.5): Approachability, Age, and Self-similarity. Approachability (see Table 2 for detailed descriptions) was the most prevalent theme among ethnic groups, with participants indicating to perceive a welcoming or friendly look from the male appearance they selected: e.g., “The counselor looked friendly which will help me to share things without hesitation” [MA028], “The only explanation I can give is that the person I selected that the most empathetic and caring look to them” [MA208]. See Appendix A, Table 12 for details on the thematic analysis per ethnic group for male appearances’ selection.
Next to Approachability, Age prevailed in all ethnic groups but the Native American/Alaska Native one: it could have been due to the small number of participants from that group (n=2). Also, Self-similarity was a top three theme among our Asian (n=217), Afro/Black American (n=28), Latin American (n=100), and Pacific Islander (n=8) participants. It is important to note that, out of the n=262 responses labeled under Age, n=70 also referred to age-matching between the virtual human and the user, comprising 58.8% of responses labeled with Self-similarity: e.g., “Looks closer to my age; can relate to people that are my age. Attaching a specific archetype to a virtual counselor is difficult” [MA125], “I feel like that virtual counselor is more close to my age which would help them relate to what I’m going through” [MA280].
These findings suggest that the agent’s age plays a crucial role in designing virtual humans tailored to assist computer science students with their mental health needs. Moreover, these findings indicate that user-agent similarity is an important factor for encouraging computer science students to engage in mental health discussions with male virtual humans.
3.7. Findings: Preferences on encouraging virtual human female appearances
Participants (n=481) were asked to choose up to three female appearances based on how they could encourage someone similar to them to engage in mental wellness conversations. Subsequently, they selected their top female appearance and answered two questions regarding their perceptions of the agent’s age and ethnicity. Figure 5 displays the four female appearances most frequently included in participants’ top three selections, as well as the four least selected female appearances. Figure 6 presents the appearances that participants most often chose as their top choice. Three of the four appearances from Figure 5 were also mostly preferred as top one choices, as shown in Figure 6. Figure 6 features a new female appearance (F7). She was the seventh most selected in participants’ top three selections of female appearances, with 21.4% of participants selecting her.
Figure 5. Most and Least Selected Female Male Appearances
3.7.1. Perceptions of age and ethnicity of virtual human female appearances
Using the Pearson correlation coefficient, we found that participants’ self-reported ages (n=479, as two participants did not report their age) significantly correlated with the ages of the female agents they selected as their top-one choice ($r=0.187,n=479,p<0.01$). This positive correlation suggests the presence of user-agent age similarity when participants selected the appearances of their female virtual agents.
Participants’ perceptions of the four top female appearances’ ages and ethnic groups are shown in Table 4. As observed in Table 4, the top agents were between 26 and 35 years of age, with the top three (F1, F2, and F3) close to the general average age of 28.6 years previously reported (Section 3.4).
Ethnic groups were moderately alike among participants and their female virtual agents, with a Jaccard index of 0.33 suggesting that more than 33% of the ethnic categories were shared between them (Luaces et al., 2012). This index was lower than the index calculated for male appearance selection (0.44).
Concerning the ethnicity of the female appearances, two of the top four female appearances (Figure 6) were perceived as White (non-Latin American)–F2 and F3, while F1 was perceived as Latin American and F7 as Asian. Analyzing participants’ preferences based on their ethnic groups, it was observed that ethnic matching was not a tendency as it was for male appearances. Although 82.1% of the Afro/Black American population (n=28) selected a non-White female appearance, only a 32.1% of them preferred an ethnic-matching female appearance. Similarly, 42.6% of White (non-Latin American) participants (n=202), 14.3% of Middle Eastern/North African participants (n=14), and 12.5% of Pacific Islander participants (n=8) selected a female appearance that matched them ethnically.
Comparing these findings to those from Section 3.6.1, we note that the gender of the virtual human affected how much participants desired the agent to match them ethnically. This could be due to participants’ general preference toward female agents (Section 3.4). These findings reveal an interesting interplay between ethnicity and gender that should be addressed with further research.
Figure 6. Top Preferences of Female Appearances
Table 4. Participants’ Perceptions of Ethnicity and Age for Their Top Female Appearances
Perceived Age Perceived Ethnicity
Agent
Average
[years of age]
Standard Deviation
[years of age]
Group
% of responses
out of n
F1 (n=159) 27.2 3.3 Latin American 93.1%
F2 (n=70) 27.1 3.0 White 82.8%
F3 (n=70) 26.7 2.8 White 91.9%
F7 (n=67) 34.2 9.4 Asian 86.1%
Among the other ethnic groups, Latin American participants (n=100) were those who mostly leaned toward ethnic similarity with their selected female appearances (77.0%), followed by Asian participants (n=217): 55.3% of them selected an Asian-looking female agent. Also, both Native American/Alaska Native participants (n=2) preferred a female appearance that matched them ethnically–although further research with a broader sample population is needed concerning this ethnic group. These findings suggest that the interplay of gender and ethnicity may affect user-agent expectations concerning ethnicity. Further research is needed to understand this phenomenon fully.
3.7.2. Participants’ rationale on female appearance selection
As it was for participants’ preferences on male appearances (Table 12), the Approachability theme prevailed as the top criterion on participants’ selections of female appearances. Table 13 shows Approachability (see Table 2 for detailed descriptions) being the most prevalent theme among ethnic groups, with participants indicating to perceive a welcoming or friendly look from the female appearance they selected: e.g., “The selected virtual counselor appears to be more approachable as she looks like someone who may be able to empathize with younger age groups and has an overall welcoming look to herself” [FA038], “This counselor also appears to be laid-back, wanting to listen to my problems and help me improve” [FA476].
Different from the rationale on the selection of male appearances (see Appendix A–Table 12), Self-similarity did not showcase in the top three themes (see Appendix A–Table 13). Only the Latin American population (n=100) had it in third place, with 25.0% of them referring to it in their responses. Nevertheless, age still was the second most relevant theme among ethnic groups, with a total of 254 responses (52.8% of participants) including it as relevant: e.g., “I look for age in a counselor, someone who is more experienced, so that is why I chose that counselor” [FA106], “she looked old enough to have been through some stuff so she could talk me through something but also not too old and had an inviting look” [FA459].
Out of the responses referring to age (n=254), 21.7% of them (55 responses) referred to Self-similarity: this, compared to the 58.8% reported for male appearance selection. While not as prevalent as with male selections, some participants expressed a preference for engaging with a female agent whose age was closer to their own: e.g., “They look young which mean they would understand the things I might have to say” [FA224], “She looks youngest among the counselors. So, it is easy to approach and share our problems with the people who are in the same age group” [FA048].
These findings suggest that age is an essential factor when designing virtual humans to assist computer science students in mental health conversations, regardless of the preferred gender or ethnicity of the agent. However, participants’ preferences on gender do seem to affect user-agent ethnic similarity, as Self-similarity was less prevalent among participants’ responses concerning female appearance selections. As reported earlier (Section 3.4, and Table 4), female virtual humans were generally preferred among ethnic groups. The selection of male virtual humans could have triggered participants to try to resemble agents ethnically; however, it was not the case for female agents. Further research is needed to better understand the interplay between gender and ethnicity concerning computer science students’ gender preferences. These findings showcase an exciting phenomenon that may help virtual human designers better tailor agents in sensitive contexts such as mental health.
3.8. Findings: Participants’ rationale on virtual human voice selection
Participants (n=481) were asked to provide information about their design decisions concerning virtual human voice selections. Table 5 displays the seven themes identified through the thematic analysis conducted on participants’ criteria.
Participants’ responses were assigned multiple labels based on the themes presented in Table 5. Inter-rater reliability was assessed separately for male and female voice selections. To achieve this, two independent researchers coded 20% of the responses (Gwet, 2008). Following this, Cohen’s kappa (Gwet, 2008) (for each theme) and Jaccard indexes (Luaces et al., 2012) were calculated for both voice selections: male and female. Inter-rater reliability was high for each theme (see Appendix A–Table 14), with Cohen’s kappa scores exceeding 0.80 (Viera et al., 2005). Additionally, multi-label inter-rater reliability was also high among raters, with Jaccard indexes surpassing 0.70 in both cases (Luaces et al., 2012).
Table 5. Themes for Voice Selection
ID Theme Description
V1 Clarity
The participant referred
to the clarity of the voice.
V2 Comfort
The participant indicated finding
the voice soothing, friendly,
and/or comforting.
V3 Ethnicity
The participant considered the ethnicity,
race, or perceived origin of the voice
for their selection.
V4 Familiarity
The participant expressed finding the
voice familiar to voices they were
used to listen.
V5 Matching
The participant made their selection
because the voice fit or matched the
appearance they previously selected
for the agent.
V6 Random
The participant did not elaborate on
their rationale or provided a
nonsensical response.
V7 Self-similarity
The participant expressed finding the
voice similar to theirs.
3.9. Findings: Preferences on encouraging virtual human male voices
Among participants’ selections (n=481) on their top four male voices (see Appendix A–Table 15), the two most selected voices were North-American: 62.0% of participants selected Canada (CA_M), and 55.9% did so for United States (US_M). These two voices are closely followed by the United Kingdom (GB_M–55.1%) and Australia (AU_M–49.9%) voices. North-American voices were still in the first two places concerning participants’ top one male voice choices: 23.3% of participants preferred the male voice from the United States (US_M) while 21.2% did so for the Canadian voices (CA_M). However, instead of the British voice (GB_M–9.1%), the Philippine (PH_M–14.6%) voice was third among the most preferred male voices. Finally, the Australian voice (AU_M–12.1%) was the fourth most preferred male voice.
Similar to the findings reported by Feijóo-García et al. (Feijóo-García et al., 2023), there were voice preferences among Indian participants (n=57) based on user-agent matching regarding the place of origin. It was observed in participants’ responses when selecting the four most encouraging male voices: 63.2% of Indian participants (n=57) selected the male Indian voice (IN_M). However, this effect was not observed in their top one preference nor among participants from other places of origin.
Among the themes described in Table 5, participants (n=481) generally indicated Comfort (50.9% of responses), Matching (45.9% of responses), and Ethnicity (25.6% of responses) as their top three criteria to select male voices: e.g., “Out of the voices I selected the Canada voice best fits my counselor. It sounds the least choppy out of the selection and is a bit deep as I imagine the vc having just by looking at the image. It also sounds confident and soothing which translates to non intimidating which I feel that other students would agree that would be an attraction point for a vc” [MV009], ‘‘I chose the voice because I think it suits the picture better and I can clearly understand what they are trying to say” [MV420].
Regardless of the catalog of male voices (see Appendix A–Table 15), there were no accent similarity expectations among international participants (n=79), as 53.2% of them preferred a North-American voice: American (US_M–31.6%) or Canadian (CA_M–21.5%), and their third most preferred voice was the British one (GB_M–17.7%). Considering the themes identified in Table 5, international participants (n=79) made their choices due to agent-voice Matching (44.3% of responses), and on voice Clarity (32.9% of responses) and Comfort (49.4%). Although 20.3% of the international population (n=79) referred to the voice’s Ethnicity, Self-similarity was not a prevalent criterion among them–8.9% of them referred to it.
Self-similarity neither was a prevalent criterion among the domestic population (n=402): Only 7.2% of them referred to it. Although Comfort (51.2% of responses) and Matching (46.3% of responses) were still two of the themes mostly considered, Clarity (22.6% of responses) did not make it to the top three of themes. Instead, Ethnicity appeared to be more relevant (26.6% of responses) as participants referred to the ethnicity, race, or perceived place of origin of the voice to be matched with their preferred male appearance: e.g., “It matches his appearance and fits his nationality in the most natural way” [MV068].
These findings add to those reported by Feijóo-García et al., (Feijóo-García et al., 2023), as it was found how male appearances affected the selection of male voices in virtual human design. As seen in Table 15, participants’ choices varied when asked to match the male appearance they previously preferred. For instance, the 55.1% of preference toward the British voice (GB_M) became a 9.1% when participants selected the most preferred male voice. In general, appearance-voice matching was an essential criterion when selecting the male voice, as voice mismatches could jeopardize a student’s experience with a male virtual human (Feijóo-García et al., 2023).
3.10. Findings: Preferences on encouraging virtual human female voices
Among participants’ selections (n=481) on their top four female voices (see Appendix A–Table 16), the two most selected female voices were again North American: 60.5% of participants selected the American voice (US_F), and 52.2% did so for the Canadian one (CA_F). In contrast to the male voice selections (Section 3.9), the Irish voice (IE_F) emerged as the third most popular choice (49.9% of participants), followed by the British voice (GB_F) with 41.4% of participants favoring it. Concerning the top female voice selections, only the American (27.2% of participants) and Irish (18.7% of participants) made it to the first three places. Both were followed by the Philippine voice (PH_F–15.6of participants). Differently from male voice selections, neither British (GB_F) nor Canadian (CA_F) voices were preferred for female agents.
Similar to observations on participants’ choices of male voices (Section 3.9), Indian participants (n=57) leaned toward female voices that matched their place of origin when selecting the four most encouraging female voices: 80.7% of Indians (n=57) selected the female Indian voice (IN_F). Similarly, participants from the United States (n=354) generally included the American voice (US_F) in their four most encouraging female voices (56.2% of responses). Nevertheless, as with male voice selections, this effect was not observed in participants’ final top preferences for female voices.
Among the themes described in Table 5, participants (n=481) generally indicated Comfort (45.9% of responses), Matching (40.1% of responses), and Ethnicity (30.8% of responses) as their top three criteria to select female voices. As observed among international participants (n=79) on their preferences of male voices (Section 3.9), Self-similarity was not a prevalent criterion for their selection of female voices: only 6.3% of responses included it. On the contrary, Comfort (48.1% of responses) and Matching (43.0% of responses) were the top reasons behind participants selections: e.g., “I believe this voice has a calming effect and it fits accordingly to the persona of the counselor i chose. Anything else will probably make me want to go to a different counselor” [FV026], “it looks perfect to her and also it will helps in motivation” [FV046].
Accent similarity expectations were not observed among international participants (n=79). Among participants from this group, 36.7% preferred the American female voice (US_F), while 20.3% favored the Irish (IE_F) voice. The Indian (IN_F) voice ranked as the third most popular choice, with 17.7% of international participants preferring it. This could be attributed to the presence of Indian participants (n=52) in this group, where 26.9% preferred the Indian female voice (IN_F) and 36.7% chose the American voice (US_F). Although Ethnicity made it in the top three main themes (21.5% of responses) for female voice selections among international participants (n=79), Self-Similarity was their least relevant theme: 6.3% referred to it. Hence, user-agent voice similarity was not observed and did not emerge as a general trend among international students (n=79) when selecting female voices.
Self-similarity was also not a prevalent criterion among domestic participants (n=402) in the selection of female voices: only 9.0% of them referred to it. Instead, Comfort (45.5% of responses) and Matching (39.6% of responses) were the two themes mostly considered, followed by Ethnicity (32.6% of responses). These themes were predominant also in their selection of male voices (Section 3.9): e.g., “It sounded genuine and friendly and went well with the image selected” [FV164], “she looks like American, so that I thought that voice suits her and also the grace and elegance of the voice suits her” [FV045].
These findings contribute those presented by Feijóo-García et al. (Feijóo-García et al., 2023), concerning the interplay between female appearances and the selection of female voices in virtual human design. As presented in Table 16, participants’ choices diverged when asked to match their previously preferred female appearance with a female voice. This is observed in the decision-making of participants from India (n=57) and the United States (n=354). Although selecting their regional voices in their top four most encouraging female voices (expressing user-agent similarity expectations), they ultimately chose the female voice that best matched their preferred female appearance–no user-agent similarity was observed.
4. Study Design, Part 2: User-Designer and User-Agent Similarity in the Promotion of Mental Health Intentions
The second part of this user study followed a structure that addressed the following research questions:
  • RQ-“Designer Cues”: Which characteristics of a virtual human’s designer can be identified by computer science students during their interactions with the agent?
  • RQ-“Designer’s Legacy Effectiveness”: How effective are virtual humans in promoting computer science students’ intentions toward mental wellness conversations when co-designed by computer science students with similar demographic characteristics to their end users?
This part explored user-designer demographic similarity in a scenario where users interacted with the agents they designed. This was done to increase user-designer demographic similarity as much as possible. At least four days were given between the first and second parts of the user study to prevent participant biases concerning their own designs. Also, the agent-designer similarity was assessed as participants were asked to guess the demographic characteristics of the designer behind their agent (see Figure 1 for reference).
4.1. Participants
The second part of this user study reached out to the same adult population of graduate and undergraduate computer science students from the University of Florida (n=481). However, only 240 participants completed the second part of this user study. Participants were enrolled in the Spring of 2023. They came from a diverse range of courses: two graduate-level Human-Centered Computing courses, as well as six undergraduate courses, including two freshman, one sophomore, and three senior CS courses.
Participants’ reported ages were: 18-20 years of age (n=161), 21-25 years of age (n=69), 26-30 years of age (n=5), and over 30 years of age (n=5). Participants also reported the number of languages they were familiar with–monolingual (n=101) or multilingual (n=139), and their place of origin–country and city. Participants reported being from the United States (n=164) or other places (n=76): India (n=36), China (n=5), Venezuela (n=5), and from other countries (n=16). Participants indicated being domestic (n=196) or international (n=44) students.
Table 6. Ethnic Groups per Gender
Ethnic Groups per Gender
Ethnic Group Total Gender
Female Male Other
Asian n=107 n=44 n=61 n=2
Afro/Black American n=11 n=4 n=6 n=1
Latin American n=55 n=17 n=36 n=2
Middle Eastern/North-African n=7 n=2 n=4 n=1
Native American/Alaska Native n=1 n=0 n=0 n=1
Pacific Islander n=1 n=0 n=0 n=1
White (non-Latin American) n=90 n=32 n=52 n=6
Participants were also asked to report their gender and ethnicity for the second part of this user study. The gender distribution among participants included Female (n=88), Male (n=144), and Other (n=8)—e.g., gender-queer, non-binary, transgender. Table 6 presents the participants’ ethnic groups categorized by their gender. Each row should be considered independently, as there was a group of multi-ethnic participants (n=66) who selected more than one ethnic group.
Figure 7. Part 2: Procedure
4.2. Procedure
The second part of this user study was conducted online, asynchronously, and was between-subjects. Participants did not take more than 30 minutes to complete it. The second part happened at least four days after participants completed the first part of the user study (Section 3).
4.2.1. Experimental groups
First, participants were randomly assigned to one of three groups, each introducing one virtual human designed to talk about gratitude journaling (Flinchbaugh et al., 2012). Gratitude journaling has been a mental health topic addressed in several studies concerning virtual humans design for mental well-being support (Feijóo-García et al., 2023, 2021; Zalake et al., 2021a). For the following two groups, each participant interacted with the preferred of the two virtual humans (male or female) they designed in the first part (Section 3). This was done to maximize the similarity between the participant and the designer behind the virtual human–they were the designers. The two intervention groups were as follows:
  • Intervention group - “Explicit legacy:” The participant interacted with a virtual human designed by a computer science student whose demographic characteristics were similar to those of the participant. The virtual human explicitly described the demographic characteristics of the designer who created it: gender, age, place of origin, and ethnicity.
  • Intervention group - “Implicit legacy:” The participant interacted with a virtual human designed by a fellow computer science student whose demographic characteristics were similar to those of the participant. The virtual human did not describe the designer who created it.
The third group was used as a control group, featuring a female virtual human whose characteristics were determined based on findings from work by Feijóo-García et al. (Feijóo-García et al., 2023), and participants’ general preferences during the co-design process in part one (Section 3.4).
  • Control Group -“No legacy:” The participant interacted with a generic virtual human not designed by a fellow computer science student and that responds to previous literature and previous findings on virtual human design.
The virtual human was female because participants generally preferred a female agent: these preferences match what previous literature on virtual human genders has reported (Baylor, 2009; Feijóo-García et al., 2021; ter Stal et al., 2020). Based on Feijóo-García et al. (Feijóo-García et al., 2023), the virtual human featured the female agent ranked first in their research and a British-accented voice: the top preferred female voice from their research. Participants in the control group did not select any British-accented voice during part one (Section 3), nor selected the female appearance they interacted with in this second part (Figure 8). Therefore, the control group featured a virtual human whose appearance and voice have been reported to be engaging and whose gender responds to good practices on virtual human design. This is to compare the effectiveness of the designer’s legacy in scenarios where participants interact with virtual humans designed by someone who shares their own demographics.
Figure 8. Control Group - “No Legacy”: Virtual Human Appearance
4.2.2. Virtual humans generation
A total of 154 virtual humans were created among the experimental groups: one for the control group and 153 for both intervention groups (Section 4.2.1). To do so, the artificial intelligence (AI) video generation platform Synthesia (Synthesia, 2022) was used in addition to its application programming interface (API) (syn, 2023) for Python. For both intervention groups, appearances and voices were gathered from participants’ input during the first part of the user study (Section 3). Figure 9 illustrates the virtual human generation process.
Figure 9. Virtual Humans Generation
The message among groups was alike. For the Control Group -“No legacy” and the Intervention group - “Implicit legacy”, the message fully referred to gratitude journaling (Flinchbaugh et al., 2012). For the Intervention group - “Explicit legacy”, the message started by indicating that the virtual human was designed by someone similar to the participant and then proceeded to disclose the designer’s age, gender, ethnic groups, and place of origin: e.g., “Hello and welcome. I am a mental wellness virtual counselor at the University of Florida. I was designed by a fellow college student who shares similar characteristics to yours. He reported being Multiracial; with Latino and White roots, and is also from Vera. He is 20 years old. I want to share with you the power of gratitude in promoting mental health and life satisfaction. Gratitude as a…”. The demographics of the designer were provided at the start of the interaction to ensure that the majority of participants in that group had access to this information before delving into the topic. Appendix C presents a sample of a script for an agent from the explicit group. Appendix B presents the script agents used in the control and implicit groups.
The gender of the designer was disclosed through pronouns, using He for participants self-identified as Male, She for those self-identified as Female, and They for those who self-identified in other gender groups. The age was first rounded to the closest multiple of five and disclosed directly in the message: e.g., for an 33 years of age, the agent disclosed an age of 35.
4.2.3. Intervention
Participants were asked to respond to an online questionnaire, in which they were assessed on their intentions toward gratitude journaling (Flinchbaugh et al., 2012) before and after interacting with a virtual human co-designed in the first part of the user study (Section 3). Participants were also asked to describe and explain their description of the designer behind the virtual human they interacted with, responding to questions concerning the designer’s place of origin, ethnicity, age, and gender. Questions were organized in five blocks (Figure 7 illustrates the procedure):
Block 1-Informed Consent provided an informed consent document with details of the study and two multiple-choice questions. Participants did not have to respond to any questions here, as they already gave their consent in the first part of the user study (Section 3).
Block 2 - Pre-TPB presented seven 7-point semantic differential scales based on the theory of planned behavior (TPB) (Zalake et al., 2021a; Ajzen, 1985) as a pre-interaction assessment. All scales were continuous (included decimal numbers). These scales align with those previously outlined by Feijóo-García et al. (Feijóo-García et al., 2023), originally introduced by Zalake et al. (Zalake et al., 2021a).
Block 3 - Virtual Human Interaction introduced a virtual human to talk about gratitude journaling (Flinchbaugh et al., 2012). The interaction lasted no more than four minutes.
Block 4 - Post-TPB presented the same seven 7-point semantic differential scales from block 2 as a post-interaction assessment.
Block 5 - Participant’s Perceptions of The Designer’s Demographics asked participants to describe the designer who created the virtual human they interacted with. The block included five questions concerning the designer’s demographic characteristics: age (open-ended numeric input), ethnicity (multiple-choice close-ended question based on (Feijóo-García et al., 2021, 2022)), place of origin (open-ended input as country-city), and gender (close-ended multiple-choice question based on (CSUSM, 2022)). The final open-ended question gathered participants’ insights about their selections.
Block 6 - Participant Demographics focused on participants’ demographics, presenting six questions. First, it asked about participants’ age (open-ended numeric input), ethnicity (multiple-choice close-ended question based on (Feijóo-García et al., 2021)), and gender (close-ended multiple-choice question based on (CSUSM, 2022)). Then, it followed by asking about participants’ international student status (domestic–United States, or international) and their place of origin (open-ended input as country-city). Finally, it asked one open-ended question to have participants list all languages used to communicate.
4.3. Data analysis
Participants’ responses were first analyzed concerning how they described the designers behind the virtual humans they interacted with. Responses were studied using descriptive statistics crossing participants’ responses with their demographic data and the actual demographics of the designers behind the virtual humans. The demographics considered were gender, ethnicity, age, and place of origin.
Participants’ rationale was qualitatively analyzed by coding participants’ responses from the open-ended question on participants’ insights about their selections to determine if they selected the designers’ demographics based on the demographics of the virtual humans. Two classes were used for this process: Similarity and Non-Similarity. Cohen’s kappa was used to calculate inter-rater reliability (Gwet, 2008; Viera et al., 2005).
Then, the analysis centered on participants’ intentions on doing gratitude journaling (Flinchbaugh et al., 2012), following the intervention previously described (Section 4.2.3). The analysis considered participants’ scores on the 7-point semantic differential scale Intention: “I intend to write down what I feel grateful about for the next week to have a good mental health” (1-false, 7-true).
(1) $Score_{N}=\begin{cases}(post-pre)/(7-pre)&post>pre\\ (post-pre)/pre&post\leq pre\end{cases}$
Participants’ responses were normalized based on their pre-interaction and post-interaction scores (Equation 1) (Zalake et al., 2021a; Feijóo-García et al., 2023). Due to the non-Gaussian distribution of responses, the non-parametric Kruskal-Wallis Test (Kruskal and Wallis, 1952; McKight and Najab, 2010) was used to compare the three experimental groups and evaluate the following hypothesis:
  • H-“User-Designer Similarity Effect”: Virtual humans co-designed by computer science students are more effective in promoting these students’ intentions toward gratitude journaling when there is a higher degree of demographic similarity between the user and the designer and the virtual human explicitly communicates the designer’s demographics.
4.4. Findings: Participants’ perceptions and rationale on designers’ demographics
Participants from the Intervention Group - “Implicit Legacy” (n=75) were assessed on their capacity to accurately tell the demographic characteristics of the designer behind the virtual human they interacted with. These virtual humans were initially designed by themselves in the first part of the user study (Section 3) and were intended to support students similar to them in mental health conversations.
Out of the four demographic characteristics of the designer, age was the characteristic that participants could better identify: 56.0% of them correctly identified the designer’s age. Two demographic characteristics followed, with 37.3% of them correctly identifying gender, and 32.0% ethnicity. The designer’s country of origin was only identified by 21.3% of participants.
Participants from the Intervention Group - “Explicit Legacy” (n=78) were also assessed concerning the designer’s demographic characteristics. For this group, the virtual human explicitly mentioned the designer’s characteristics at the beginning of the interaction–no longer than four minutes. Surprisingly, participants from this group did not fully recall the four characteristics initially mentioned at the beginning of the interaction (Section 4.2.2). The designer’s country of origin was the characteristics participants from this group mostly recalled (50.0% of participants). It was followed by ethnicity (44.9% of participants), age (43.9% of participants) and gender (42.3% of participants).
To better understand the interplay between the agent, the designer, and the participant, especially for the Intervention Group - “Explicit Legacy” (n=78), participants’ perceptions were analyzed concerning three of the virtual humans’ demographics they were asked to identify in the first part of the user study: age, gender, and ethnicity (Sections 3.6.1 for male appearances and 3.7.1 for female ones). Concerning gender, 75.6% of the participants from this group considered the agent’s gender to match the designer’s, while 64.1% of them did so concerning the agent’s ethnicity. Of the three demographic characteristics, age was the only one that participants did not generally relate to the designer: only 46.2% of them considered the agent’s age to match the designer’s. Looking toward the participants from the Intervention Group - “Implicit Legacy” (n=75), 80.0% of them considered the agent’s gender to match the designer’s, and 78.7% of them did so with the agent’s ethnicity. Also, although lower, over half of that group related the agent’s perceived age with the designer’s (56.0% of participants).
Participants from the Control Group -“No legacy” (n=87) also guessed who was the designer behind the virtual human. Considering their perceptions of the agent’s demographics, it was observed that 64.4% of them related the agent’s gender to the designer’s, while 59.8% did so for ethnicity. Also, 62.1% of participants related the agent’s and designer’s age.
Participants (n=240) were also asked to share their insights behind their responses to the four demographic characteristics they were asked to identify: gender, age, ethnicity, and place of origin.
Participants’ responses were coded into two classes to label if they used or not agent-designer similarity to determine the designer’s demographic characteristics. Then, Inter-rater reliability was used to categorize participants’ open-ended responses. For this, an independent researcher coded 20% of the responses (Gwet, 2008). Following, another independent rater coded that sub-set of responses. Cohen’s kappa was calculated to measure agreement among raters, getting a score of 0.81– a high level of agreement between raters (Viera et al., 2005).
In general, participants from the three experimental groups (Section 4.2.1) considered agent-designer similarity to identify who was behind the virtual human. As observed in Figure 10, both the Intervention Group - “Implicit Legacy” and the Control Group - “No Legacy” had over 75% of their participants referring to agent-designer similarity. Close to them, 66.7% of the participants from the Intervention Group - “Explicit Legacy” also considered agent-designer similarity when determining the designer’s demographic characteristics: e.g., “From her appearance, I could clearly figure out that she is in her 40s and belong to south-asian origin” [DR028], “Ethnic group and gender were based on how they look. Place of origin is because it was in front of a UF background. Age was a guest, I’m not sure who is conducting this research, but I suppose maybe on the younger side” [DR213], ‘‘She could have been born anywhere I guess but based on her looks her ethnicity seemed to be South Asian so I said India. I didn’t know what part so I just thought of the first city that came to mind. I was more confident about her gender and age, and I guessed 27 because she seemed young” [DR073].
Figure 10. Participants’ Rationale on Agent-Designer Similarity
These findings reveal important points about the connection between the demographic characteristics of virtual humans and those of their designers. First, participants tend to relate the demographics of virtual humans to those of their designers–designer-agent demographic similarity, particularly in regard to ethnicity and gender. Considering how participants were co-designers in the first part of the study (Section 3), these findings suggest that computer science students acting as designers could be influenced by their own demographics when creating virtual humans.
Second, these findings show that the mere awareness of the designer’s demographic characteristics is insufficient for a participant to accurately identify who created the virtual human. This highlights the limitations of demographic information in determining the creators of virtual human designs.
Finally, it was observed that the visual cues of a virtual human could overshadow verbally shared information in their message, as seen with the Intervention Group - “Explicit Legacy” (n=78) who had difficulty recalling the designer’s demographics initially provided. This finding implies that the visual aspects of virtual human demographics could have a stronger impact on users’ perceptions than verbal cues, especially when agent-designer demographic similarity is present in participants’ rationale when determining the designer’s characteristics.
4.5. Findings: Designer cues and participants’ intentions towards gratitude journaling
Following the strategy presented in Section 4.3, the three experimental groups (Section 4.2.1) were compared in regard to participants’ intentions towards gratitude journaling (Flinchbaugh et al., 2012).
Table 7. Descriptive Statistics - Normalized TPB Score of Intention
Experimental Group N Mean SD SE Coefficient of variation
Explicit Legacy $78$ $0.236$ $0.37$ $0.04$ $1.65$
Implicit Legacy $75$ $0.20$ $0.34$ $0.04$ $1.70$
No Legacy (Control) $87$ $0.24$ $0.36$ $0.04$ $1.50$
Experimental groups were compared based on participants’ normalized scores of intentions (Equation 1). The non-parametric Kruskal-Wallis (Kruskal and Wallis, 1952; McKight and Najab, 2010) was used to validate the following hypothesis H-“User-Designer Similarity Effect”: Virtual humans co-designed by computer science students are more effective in promoting these students’ intentions toward gratitude journaling when there is a higher degree of demographic similarity between the user and the designer and the virtual human explicitly communicates the designer’s demographics. The hypothesis was rejected, as no significant difference was observed among groups concerning participants’ normalized scores of intention (Kruskal-Wallis Test, H(2) = 1.06, p=0.59). Figure 11 shows that the three groups presented a similar distribution with no visible differences among them (see Table 7 for detailed descriptive statistics), with normalized scores generally over 0.0.
Figure 11. Normalized Intentions per Experimental Group
Experimental groups were also compared concerning participants’ post-TPB scores on intention (Section 4.2.3. No significant difference was observed among groups concerning post-TPB scores (Kruskal-Wallis Test, H(2) = 0.29, p=0.86).
Figure 12. Post-TPB Intentions per Experimental Group
Just as with participants’ normalized scores, the post-TPB scores were similarly distributed among groups (Figure 12) with means around 4.0 on a 7-point scale (Table 8).
Table 8. Descriptives - Post-TPB Score: Intention
Experimental Group N Mean SD SE Coefficient of variation
Explicit Legacy $78$ $4.108$ $2.080$ $0.236$ $0.506$
Implicit Legacy $75$ $4.197$ $1.874$ $0.216$ $0.446$
No Legacy (Control) $87$ $4.318$ $1.887$ $0.202$ $0.437$
The lack of significance can be interpreted from several angles, considering the findings previously discussed in Section 4.4. It is possible that the demographic characteristics of the designers and their impact on the design of the virtual humans were not as influential on participants’ intentions toward gratitude journaling. The findings of this work indicate that the degree of demographic similarity between the user and the designer, and the virtual human’s explicit communication of the designer’s demographics, did not significantly differentiate the groups based on the participants’ normalized intentions scores.
Second, this work’s findings suggest that the demographic information of the designers, which the participants were aware of, did not provide sufficient cues to accurately identify the designer of the virtual human. Thus, this awareness may not have been strong enough to influence participants’ intentions toward promoting gratitude journaling.
Lastly, the visual cues of the virtual humans could have overshadowed the information verbally communicated by the virtual human regarding the designer’s demographics. The fact that the Intervention Group - “Explicit Legacy” (n=78) struggled to consistently remember the designer’s demographics initially provided could serve as an indication that the visual aspects of the virtual human’s demographics might have dominated their perceptions, affecting their intentions toward gratitude journaling.
Overall, these findings suggest that tailoring designer cues in virtual human design is a complex task. The visual characteristics of virtual humans and users’ expectations (or biases) could influence users’ intentions beyond the effects of any demographic similarities between users and designers; even with virtual humans explicitly communication their designer’s demographics.
5. Discussion and Future Work
This section is structured to discuss our findings in regard to the three research questions initially posed in Section 1:
  • RQ-“Similarity Elucidation”: How do similarity expectations influence the design decisions made by computer science students when co-designing virtual human for mental health support?
  • RQ-“Designer Cues”: Which characteristics of a virtual human’s designer can be identified by computer science students during their interactions with the agent?
  • RQ-“Designer’s Legacy Effectiveness”: How effective are virtual humans in promoting computer science students’ intentions toward mental wellness conversations when co-designed by computer science students with similar demographic characteristics to their end users?
5.1. Insights on agent-designer similarity in virtual humans co-design
Responding to the research question RQ-“Similarity Elucidation”: How do similarity expectations influence the design decisions made by computer science students when co-designing virtual humans for mental health support?, this work’s findings provide valuable insights into the design of virtual humans to support college students in mental health conversations, especially those majoring in computer science. From the different factors identified in Section 3.4, age was found to be an essential design factor, regardless of participants’ genders or ethnic groups, for the design of virtual humans to support mental health. Although agent-student age matching was not as predominant as other demographic characteristics (e.g., place of origin), it was observed that participants’ rationale generally considered age for both male and female appearance selection (Sections 3.6.2 and 3.7.2). Hence, the appropriate tailoring of age can help promote effective virtual human interventions in assisting computer science students (and other potential populations) in mental health conversations.
Notably, the preference for ethnic matching differed between male and female appearances (Sections 3.6.2 and 3.7.2). While ethnic matching was more prevalent in the selection of male appearances, it was less pronounced for female appearances. This suggests that computer science students may have different expectations and criteria regarding the ethnicity of male and female virtual humans. As female virtual humans were generally preferred among ethnic groups, male virtual humans may have triggered participants toward a need for user-agent ethnic similarity. These findings suggest that gender and ethnicity interact in a complex way and highlight the importance of considering user-agent similarity and their intersectionality in designing virtual humans for mental health support. More research is needed to explore these gender-specific preferences and the interplay between ethnicity and gender in virtual human design.
Regarding participants’ preferences for virtual human voices, it was observed that appearance-voice matching played a crucial role, particularly in the design of male virtual humans (Section 3.6.2). This observation was particularly observed among domestic participants, for whom ethnicity held greater significance compared to international participants. They prioritized appearance-voice matching, voice clarity, and voice comfort (Table 5). These findings support previous work by Feijóo-García et al. (Feijóo-García et al., 2023), stating the importance of appearance-voice matching in virtual humans’ effectiveness in promoting mental health conversations.
Ethnicity also emerged as a notable criterion when selecting voices (Sections 3.9 and 3.10), with some participants even expressing a preference for voices that matched their own ethnic background. It was found that user-agent ethnic similarity has a major role in encouraging computer science students to engage in mental health discussions with male virtual humans, as participants preferred male virtual humans who shared age and ethnicity. This indicates that computer science students may feel more comfortable and open when interacting with male virtual humans who are closer in age and ethnic background. Hence, correctly modeling these factors may help in designing more effective virtual humans.
This work’s findings have significant implications for designing virtual humans to support college students in sensitive contexts such as mental health. Age, gender, ethnicity, appearance-voice matching, and user-agent similarity are essential factors. Virtual human designers should aim to understand the preferences and expectations of their target population (e.g., computer science students) when creating virtual humans. Correctly tuning characteristics and agent-user similarity can determine the success of virtual human interactions.
It is important to acknowledge the limitations of this work, as the findings are derived from a specific sample of participants in one major North American college. Future research should aim to replicate and expand upon these findings with more extensive and diverse sample populations to understand better the complex interplay between age, gender, ethnicity, appearance, and voice in virtual human design for mental health contexts.
These findings contribute to understanding factors influencing computer science students’ preferences and expectations regarding virtual humans in mental health conversations. The insights gained from this work can inform the design and development of virtual humans that effectively support and engage students in meaningful mental health discussions.
5.2. Insights on the designer’s legacy and users’ perceptions of it in virtual human interactions
Responding to the research question RQ-“Designer Cues”: Which characteristics of a virtual human’s designer can be identified by computer science students during their interactions with the agent?, this work’s findings provide valuable insights into the connection between the demographic characteristics of virtual humans and those of their designers (Section 4.4). One key finding is that participants tended to associate virtual humans’ demographics with their designers’ demographics, particularly regarding ethnicity and gender. Although relevant in the selection of virtual human appearances (Sections 3.6.2 and 3.7.2), the designer’s age was not identifiable nor recognizable through the virtual human. Same for the place of origin: the characteristic that most likely presented a similarity-attraction effect (Section 3.4). Considering how participants were co-designers in the first part of the user study (Section 3), these findings suggest that computer science students acting as designers could be influenced by their own demographics when creating virtual humans.
It was observed that the mere awareness of the demographic characteristics of the virtual human’s designer was insufficient for participants to accurately identify the virtual human’s designer (Section 4.4). As reported in Section 4.4, participants’ perceptions led them to perceive the designer as demographically similar to their agent–designer-agent demographic similarity. Moreover, we observed a more intriguing phenomenon where the visual cues of the virtual human appeared to overshadow the verbally shared information regarding the designer’s demographics. Specifically, participants who interacted with a virtual human that verbally shared the demographic information of its designer had a tough time recalling the designer’s information. This finding suggests that the visual aspects of virtual human demographics hold more influence over users’ perceptions compared to verbal cues, and it may be more pronounced when participants base their rationale on agent-designer similarity. This highlights the limitations of demographic information in determining the creators of virtual human designs.
These findings contribute insights to help tailor designer cues in virtual humans for mental health support. As observed, user-designer and agent-designer similarity interact in a complex manner, even when computer science students serve as designers and users of virtual humans. Future research should explore different ways to present designer cues, adding factors apart from demographic characteristics, appearances, and voices.
5.3. Insights on the designer’s legacy effectiveness in virtual human interactions
Responding to the research question RQ-“Designer’s Legacy Effectiveness”: How effective are virtual humans in promoting computer science students’ intentions toward mental wellness conversations when co-designed by computer science students with similar demographic characteristics to their end users?, it was found that virtual humans promoting computer science students’ intentions toward mental wellness conversations are not more effective when co-designed by fellow computer science students with similar demographic characteristics to their end users (Section 4.5).
We observed that the visual cues of the virtual humans could have overshadowed the information verbally communicated by the virtual human regarding the designer’s demographics. The fact that the Intervention Group - ”Explicit Legacy” (Section4.2.1) struggled to recall the designer’s demographics initially provided by the virtual human could serve as an indication that the visual aspects of the virtual human’s demographics might have dominated their perceptions, potentially affecting their intentions toward gratitude journaling. It was also observed that the demographic information of the designers, which the participants were aware of, did not provide sufficient cues to accurately identify the virtual human’s designer. Thus, this awareness could not have been strong enough to influence participants’ intentions toward promoting gratitude journaling.
Overall, our research suggests that tailoring designer cues in virtual human design is a complex task. This, as the visual aspects of virtual humans and users’ biases could influence users’ intentions regardless of 1) the demographic similarity between the user and the designer and 2) the virtual human explicitly communicating the designer’s demographics. Future research is needed to understand how to tailor designer cues and assess their impact on promoting intentions and behaviors toward mental health conversations.
6. Limitations
Several limitations should be acknowledged when interpreting the findings from this study. Initially, participants used the artificial intelligence (AI) video generation platform Synthesia (Synthesia, 2022) and its application programming interface (API) (syn, 2023) to create virtual humans. Although the platform provided a diverse catalog of female and male agents, the choices were restricted to the designs available in the catalog. Notably, this catalog only featured male and female agents, which did not accommodate non-binary gender expressions. Furthermore, participants, particularly those from Latin American backgrounds, found the options for male appearances limited. Future research should aim to expand the design capabilities of virtual humans, allowing for a more nuanced interplay of gender, race, ethnicity, and age to better reflect the voice and appearance of the virtual agents.
In addition to the limitations posed by the platform’s catalog, the first part of our user study also faced challenges with the diversity and balance of its sample population. Although the number of participants was increased from previous efforts (Feijóo-García et al., 2023) and had a larger population of female participants (n=182), the group remained predominantly male (n=281) and also lacked sufficient representation from some ethnic backgrounds. Our user study depended on the student population we could recruit from our institution at the time and was limited to its socio-geographic context. We acknowledge that we could have recruited a more balanced demographic sample through platforms that allow demographic recruitment specifications (e.g., Prolific (Palan and Schitter, 2018)). However, we were limited by our resources and relied on a sample population (n=481) originating from the University of Florida. Also, we acknowledge that the decision to include only computer science students could limit the broader applicability of the findings, as it does not represent a cross-section of the general population. Please keep in mind that we limited our population to this context as previous literature suggests computer science students experience the highest rates of anxiety and depression among STEM college students (Danowitz and Beddoes, 2018), with female computer science students being particularly more at risk (Rosenstein et al., 2020; Soares Passos et al., 2020). We recommend considering these limitations to interpret our findings and observations better.
Another possible limitation concerns the reuse of the same sample of participants across different stages of the study, which could have introduced biases, especially in the second part of the user study: Their previous experiences designing the virtual agents might have influenced their subsequent perceptions of the agents and their designers. Moreover, although we gave at least four days between the parts of the study, this period might not have been sufficient to prevent carryover effects, potentially affecting the outcomes. These limitations should be considered when interpreting our observations, and they may help explore further research in this area.
The study also did not sufficiently explore the influence of factors beyond gender and ethnicity, such as facial expressions and physical attractiveness, on participants’ perceptions of virtual humans. For example, the top-selected male virtual humans predominantly displayed happy expressions, which might have influenced participant choices and highlighted a potential oversight regarding the impact of facial cues. When interpreting our observations, we should acknowledge these limitations and consider their implications for the study’s design and outcomes. By doing so, future research can be better positioned to address these shortcomings, which may help explore further research in this area.
Moreover, in the second part of the study, it was observed that participants from the Intervention group - “Explicit legacy” (Section 4.2.1) struggled to fully recall the designer’s demographic characteristics. This suggests that future research should explore alternative methods to mitigate this issue. However, this limitation also provided valuable insights: designer cues should not be limited to verbal messages. These findings showed that the visual demographics of the agent could overshadow the verbally communicated demographics of the designer, indicating that the design process should carefully consider the balance between these two forms of communication.
These limitations provide helpful insights for future research to address the complexity involved in designing virtual humans and understanding the impact of their demographic characteristics on user perceptions.
7. Conclusions
Our work contributes insights concerning agent-user, agent-designer (which we refer as the designer’s legacy), and user-designer demographic similarity in the co-design of virtual humans, reporting on how computer science students, when acting as designers, reflect their own characteristics in their virtual humans. Moreover, our work contributes insights into how computer science students, when acting as users, expect virtual humans to resemble their designers.
We found participants preferred voices that matched their own ethnic background, suggesting that user-agent ethnic similarity could encourage more active engagement in mental health discussions. Moreover, we observed that age, gender, ethnicity, appearance-voice matching, and user-agent similarity emerged as critical factors in the design of virtual humans for mental health support.
Additionally, participants generally associated the demographic characteristics of virtual humans with those of their designers, particularly in terms of ethnicity and gender. Our findings also suggest that the visual cues of the virtual human could overshadow the verbally shared information about the designer’s demographics, indicating that visual aspects have more influence over users’ perceptions.
Finally, virtual humans were not found to be more effective in promoting mental health conversations when co-designed by computer science students with similar demographic characteristics to their end users. However, observations indicate that the visual cues of the virtual humans overshadowed the information verbally communicated about the designer’s demographics, possibly affecting participants’ intentions toward mental health conversations. Thus, the demographic information of the designers was not sufficient to influence participants’ intentions toward mental health conversations, suggesting that tailoring designer cues in virtual human design is a complex task that should be explored further, especially in sensitive scenarios such as mental health support.
Future directions should further explore the role of user-designer demographic similarity to understand how designer cues should be modeled for virtual human interactions. Also, future next steps should further investigate how user-agent and user-designer characteristics should be matched and featured in user-agent interactions and their impact on users’ intentions and behaviors toward mental health conversations.
Acknowledgements.
The authors wish to thank the students who actively participated in the user study. Additionally, we thank Dr. Sharon Chu, Dr. Laura Cruz Castro, Mr. Amanpreet Kapoor, Dr. Yingbo Ma, and Ms. Lisha Zhou for letting us recruit their students. This work was sponsored by the Fulbright Scholar Program and the Colombian Ministry of Science, Technology, and Innovation. This work was also supported by the National Science Foundation award numbers 1900961 and 1800947, the American foundation for Suicide Prevention (LSRG-1-050-18), and the National Institute of Mental Health (NIMH1R34MH119294-01).
Appendix A Part 1: Additional Tables
Table 9. Part 1: Participants’ Countries of Origin
Participants’ Countries of Origin
Country Participants Country Participants
ALBANIA 1 KOREA 2
ARGENTINA 1 MEXICO 1
BANGLADESH 1 NETHERLANDS 1
BARBADOS 1 PERU 1
BRAZIL 7 PHILIPPINES 3
CANADA 1 POLAND 1
CHINA 16 PUERTO RICO 1
COLOMBIA 2 SOUTH KOREA 1
COSTA RICA 1 SPAIN 2
CUBA 6 TAIWAN 1
DOMINICAN REPUBLIC 1 TANZANIA 1
EGYPT 1 THAILAND 1
GERMANY 1 TURKEY 1
INDIA 57 UNITED KINGDOM 1
INDONESIA 1 UNITED STATES 354
JAMAICA 1 VENEZUELA 7
JAPAN 1 VIETNAM 2
Table 10. Part 1: Participants’ Preferences on Gender Based on Ethnic Groups
Ethnic Group Total Preferred Gender
Female Male Other
Asian n=217 57.6% 35.5% 6.9%
Afro/Black American n=28 53.6% 32.1% 14.3%
Latin American n=100 56.0% 32.0% 12.0%
Middle Eastern/North African n=14 64.3% 21.4% 14.3%
Native American/Alaska Native n=2 50.0% 50.0% 0.0%
Pacific Islander n=8 62.5% 25.0% 12.5%
White (non-Latin American) n=202 45.0% 41.1% 13.9%
Table 11. Part 1: Inter-rater Reliability Scores for Female and Male Appearance Selection
ID Theme
Male Appearances
(Jaccard Index= 0.78)
Female Appearances
(Jaccard Index= 0.81)
Cohen’s kappa Cohen’s kappa
A1 Age 0.99 0.98
A2 Approachability 0.91 0.88
A3 Attire 0.92 0.98
A4 Ethnicity 0.98 1.00
A5 Expertise 0.94 0.89
A6 Gender 0.98 1.00
A7 Non-verbal Language 0.92 0.95
A8 Random 0.96 0.98
A9 Self-similarity 0.91 0.91
A10 Similarity to Trusted Agent 0.97 0.96
Table 12. Part 1: Thematic Analysis per Ethnic Group for Male Appearances’ Selection
ID Count
Asian
(n=217)
Afro/Black
American
(n=28)
Latin
American
(n=100)
Middle
Eastern/
North
African
(n=14)
Native
American/
Alaska
Native
(n=2)
Pacific
Islander
(n=8)
White
(n=202)
A1 262 54.8% 46.4% 48.0% 64.3% 0.0% 62.5% 55.9%
A2 264 48.4% 39.3% 59.0% 71.4% 50.0% 37.5% 58.9%
A3 75 15.2% 17.9% 15.0% 21.4% 0.0% 12.5% 18.3%
A4 46 16.1% 10.7% 7.0% 0.0% 0.0% 12.5% 2.5%
A5 110 20.7% 17.9% 24.0% 21.4% 100.0% 12.5% 25.7%
A6 13 1.8% 3.6% 2.0% 7.1% 0.0% 12.5% 3.5%
A7 74 15.2% 10.7% 11.0% 42.9% 0.0% 12.5% 15.8%
A8 8 0.9% 0.0% 4.0% 0.0% 0.0% 0.0% 1.5%
A9 119 27.2% 32.1% 29.0% 21.4% 0.0% 62.5% 20.8%
A10 70 14.3% 25.0% 14.0% 7.1% 50.0% 25.0% 13.4%
Table 13. Part 1: Thematic Analysis per Ethnic Group for Female Appearances’ Selection
ID Count
Asian
(n=217)
Afro/Black
American
(n=28)
Latin
American
(n=100)
Middle
Eastern/
North
African
(n=14)
Native
American/
Alaska
Native
(n=2)
Pacific
Islander
(n=8)
White
(n=202)
A1 254 51.2% 46.4% 53.0% 57.1% 0.0% 87.5% 54.0%
A2 355 61.8% 50.0% 67.0% 71.4% 100.0% 50.0% 61.4%
A3 54 8.3% 10.7% 6.0% 14.3% 0.0% 0.0% 12.4%
A4 30 8.3% 3.6% 10.0% 0.0% 0.0% 0.0% 0.5%
A5 141 27.6% 32.1% 22.0% 7.1% 50.0% 25.0% 22.8%
A6 11 2.8% 0.0% 0.0% 7.1% 0.0% 0.0% 2.0%
A7 117 21.2% 17.9% 19.0% 35.7% 0.0% 25.0% 19.8%
A8 16 1.4% 7.1% 4.0% 0.0% 0.0% 12.5% 3.0%
A9 122 20.3% 21.4% 25.0% 28.6% 0.0% 25.0% 20.3%
A10 91 14.7% 10.7% 20.0% 7.1% 50.0% 37.5% 15.3%
Table 14. Part 1: Inter-rater Reliability Scores for Voice Selection
ID Theme
Male Voices
(Jaccard Index=0.73)
Female Voices
(Jaccard Index=0.75)
Cohen’s kappa Cohen’s kappa
V1 Clarity 0.86 0.92
V2 Comfort 0.86 0.86
V3 Ethnicity 0.84 0.84
V4 Familiarity 0.97 0.95
V5 Matching 0.93 0.90
V6 Random 0.98 0.97
V7 Self-similarity 0.94 0.96
Table 15. Part 1: Male Voices Preferences
Male Voices
ID Name Accent Origin
% participants selecting
it in their top four
% participants selecting
it as their top one
AU_M AU Natural Australia 49.9% 12.1%
CA_M CA Natural Canada 62.0% 21.2%
IN_M IN Natural 4 India 19.5% 1.2%
IE_M IE Natural Ireland 32.6% 5.8%
KE_M KE Natural Kenya 21.0% 3.1%
NZ_M NZ Natural New Zealand 25.2% 4.4%
NG_M NG Natural Nigeria 19.3% 2.5%
PH_M PH Natural Philippines 33.9% 14.6%
ZA_M ZA Natural South Africa 16.4% 3.3%
TZ_M TZ Natural Tanzania 14.1% 0.6%
GB_M GB Natural United Kingdom 55.1% 9.1%
US_M US Natural United States 55.9% 23.3%
Table 16. Female Voices Preferences
Part 1: Female Voices
ID Name Accent Origin
% participants selecting
it in their top four
% participants selecting
it as their top one
AU_F AU Natural Australia 39.7% 6.0%
CA_F CA Natural Canada 52.2% 9.4%
IN_F IN Natural India 27.9% 5.6%
IE_F IE Natural Ireland 49.9% 18.7%
KE_F KE Natural Kenya 18.3% 1.7%
NZ_F NZ Natural New Zealand 24.7% 4.0%
NG_F NG Natural Nigeria 21.2% 2.3%
PH_F PH Natural Philippines 38.5% 15.6%
ZA_F ZA Natural South Africa 26.8% 3.7%
TZ_F TZ Natural Tanzania 12.3% 1.2%
GB_F GB Natural United Kingdom 41.4% 7.5%
US_F US Natural United States 60.5% 27.2%
Appendix B Part 2: Agent Script for the Explicit Group
“Hello and welcome. I am a mental wellness virtual counselor at the University of Florida. I want to share with you the power of gratitude in promoting mental health and life satisfaction.
Gratitude as a discipline helps protect us from destructive feelings such as envy, resentment, greed, and bitterness. One of the best ways to cultivate gratitude is to establish a daily practice, gratitude, journaling. When we are grateful, we affirm that sources of goodness exist in our lives. By writing each day, we magnify and expand on these sources of goodness. Setting aside time on a daily basis to recall moments of gratitude associated with even mundane or ordinary events. Personal attributes one has or valued people, one encounters has the potential to weave together a sustainable life theme of gratefulness. Just as it nourishes a fundamentally affirming life stance. Generally, gratitude, journaling needs 15 minutes per day, at least three times per week for at least two weeks.
There is no wrong way to keep a gratitude journal. But here are some general instructions that can help you get started, write down up to five things for which you feel grateful. The physical record is important, not just in your head, the things you list can be relatively small in importance. The goal of the exercise is to remember a good event, experience, person or thing in your life as you write down in the journal consider eight important tips.
First, be specific as possible. Specificity is a key to fostering gratitude. Second, go for depth over breadth, elaborate in detail about a particular person or thing for which you are grateful. Third, get personal, focusing on people to whom you are grateful has more of an impact than focusing on things for which you are grateful. Fourth try, subtraction, not addition, consider what your life would be like without certain people or things rather than just tallying up all the good stuff. See good things as gifts, try to relish and savor the gifts you have received. Fifth savor surprises, try to record events that were unexpected or surprising. Sixth revise, if you repeat writing about some of the same people and things is ok. But think aspects that appear too finally write regularly, whether you write daily or every other day, commit to a regular time to journal.
I hope you find gratitude. Journaling useful. As you may see, it is not a time consuming task and it brings with it many benefits. To your mental health.”
Appendix C Part 2: Agent Script for the Control and Implicit Groups
“Hello and welcome. I am a mental wellness virtual counselor at the University of Florida. I was designed by a fellow college student who shares similar characteristics to yours. She reported being multiracial with Latina and White roots and is also from Gainesville. She is 20 years old. I want to share with you the power of gratitude in promoting mental health and life satisfaction.
Gratitude as a discipline helps protect us from destructive feelings such as envy, resentment, greed, and bitterness. One of the best ways to cultivate gratitude is to establish a daily practice, gratitude, journaling. When we are grateful, we affirm that sources of goodness exist in our lives. By writing each day, we magnify and expand on these sources of goodness. Setting aside time on a daily basis to recall moments of gratitude associated with even mundane or ordinary events. Personal attributes one has or valued people, one encounters has the potential to weave together a sustainable life theme of gratefulness. Just as it nourishes a fundamentally affirming life stance. Generally, gratitude, journaling needs 15 minutes per day, at least three times per week for at least two weeks.
There is no wrong way to keep a gratitude journal. But here are some general instructions that can help you get started, write down up to five things for which you feel grateful. The physical record is important, not just in your head, the things you list can be relatively small in importance. The goal of the exercise is to remember a good event, experience, person or thing in your life as you write down in the journal consider eight important tips.
First, be specific as possible. Specificity is a key to fostering gratitude. Second, go for depth over breadth, elaborate in detail about a particular person or thing for which you are grateful. Third, get personal, focusing on people to whom you are grateful has more of an impact than focusing on things for which you are grateful. Fourth try, subtraction, not addition, consider what your life would be like without certain people or things rather than just tallying up all the good stuff. See good things as gifts, try to relish and savor the gifts you have received. Fifth savor surprises, try to record events that were unexpected or surprising. Sixth revise, if you repeat writing about some of the same people and things is ok. But think aspects that appear too finally write regularly, whether you write daily or every other day, commit to a regular time to journal.
I hope you find gratitude. Journaling useful. As you may see, it is not a time consuming task and it brings with it many benefits. To your mental health.”